# Better Estimation of the KL Divergence Between Language Models

### v2

@@ 3 Rao–Blackwellized Monte Carlo
See § E.1. ∎
The identity $\delta_{\textsc{mc}}^{N}=\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}$ shows that we can view $\delta_{\textsc{mc}}^{N}$ as the sum of step-wise Monte Carlo estimators $\delta_{\textsc{mc}}^{n}$. We now apply Rao–Blackwellization to each step-wise estimator $\delta_{\textsc{mc}}^{n}$. Concretely, we define
$T_{n}(\overline{\boldsymbol{Y}})=\overline{\boldsymbol{Y}}_{<n}$ (16)
and apply Rao–Blackwellization to each $\delta_{\textsc{mc}}^{n}$ as follows
$\delta_{\textsc{rb}}^{N}$ =$\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}\left[\delta_{\textsc{mc}}^{n}\mid T_{n}\right]=\sum_{n=1}^{N}\frac{1}{M}\sum_{m=1}^{M}\operatorname*{\mathbb{E}}_{\overline{Y}_{n}}\left[\log\frac{\vec{p}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}\right],$ (17)
where the expectation in Eq. 17 is the exact KL divergence between the conditional distributions and is computed as: $\sum_{\overline{Y}_{n}\in\overline{\Sigma}}\vec{p}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})\log\frac{\vec{p}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}$. The following theorem establishes that $\delta_{\textsc{rb}}^{N}$ is an unbiased estimator of $\delta_{\textsc{mc}}^{N}$, and furthermore, that $\delta_{\textsc{rb}}^{N}$ has lower or equal variance compared to $\delta_{\textsc{mc}}^{N}$.
Lemma 4.
With regard to the estimator $\delta_{\textsc{rb}}^{N}$ defined in Eq. 17, the following two properties hold
  1. 1.
    $\operatorname*{\mathbb{E}}[\delta_{\textsc{rb}}^{N}]-\operatorname*{\mathbb{E}}[\delta_{\textsc{mc}}^{N}]=0$ (unbiasedness)
  2. 2.
    $\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{rb}}^{N}}\right]\leq\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{mc}}^{N}}\right]$ (variance reduction)
@@
See § E.2. ∎
We then define the general case as $\delta_{\textsc{rb}}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\lim_{N\rightarrow\infty}\delta_{\textsc{rb}}^{N}$. The following theorem shows that $\delta_{\textsc{rb}}$ is an unbiased estimator of the KL divergence with variance less than or equal to $\delta_{\textsc{mc}}$.
Theorem 5.
Assume that both MC and RB estimators have finite variance, i.e., $\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{mc}}}\right]<\infty,\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{rb}}}\right]<\infty$, then following properties regarding $\delta_{\textsc{rb}}$ hold
  1. 1.
    $\operatorname*{\mathbb{E}}\big{[}\delta_{\textsc{rb}}\big{]}-\textsc{kl}(p\mid\mid q)=0$ (unbiasedness)
  2. 2.
    $\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{rb}}}\right]\leq\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{mc}}}\right]$ (variance reduction)
Proof.
See § E.3. ∎
Notably, $\delta_{\textsc{rb}}$ is guaranteed to be non-negative. Each step-wise estimator computes the exact KL divergence between the two distributions, conditioned on the sampled context $\boldsymbol{y}_{<n}^{(m)}$. Since all terms in Eq. 17 are non-negative, $\delta_{\textsc{rb}}$ remains non-negative as well.
Complexity analysis.
At first glance, computing the Rao–Blackwellized estimator in Eq. 17 might seem more computationally expensive than the standard Monte Carlo estimator in Eq. 6a, since it requires evaluating the exact KL divergence at each token position. However, this added computation does not change the overall runtime complexity. To sample $M$ strings of at most length $N$, we perform a forward pass through the language model for $N$ tokens, followed by the final transformation and a softmax. Assuming that the model representations live in $\mathbb{R}^{d}$, the next token prediction step requires ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MNd|\Sigma|\right)}$ operations, where $|\Sigma|$ is the size of the alphabet. Thus, the total time complexity of sampling $M$ strings is ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MT_{N}+MNd|\Sigma|\right)}$, where $T_{N}$ denotes the time complexity of performing a forward pass for a single string of length $N$. For KL estimation, computing the MC estimator requires an additional ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MN\right)}$ operations, while the Rao–Blackwellized estimator requires ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MN|\Sigma|\right)}$. Therefore, despite the additional computation in the Rao–Blackwellized estimator, its overall runtime complexity is the same as the standard Monte Carlo estimator. In general, $N$ is itself a random variable. Thus, a more fine-grained, randomized analyze would yield the bound ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MT_{\mathbb{E}[N]}+M\mathbb{E}[N]d|\Sigma|\right)}$ where $\mathbb{E}[N]$ is the expected string length under $p$.
4 Estimating the Gradient
KL estimation plays a crucial role in various applications, particularly in fine-tuning large language models to minimize KL divergence. One prominent example is reinforcement learning from human feedback (RLHF), where the objective function typically includes a KL regularization term to balance reward maximization with constraint adherence.
In this setting, the language model is a differentiable function of a real-valued parameter vector $\boldsymbol{\theta}$, and $\boldsymbol{\theta}$ is updated via gradient descent to maximize the KL-regularized RL objective. Therefore, an essential step in this optimization process is to compute the gradient of the KL divergence with respect to the model’s parameters, i.e.,
$\nabla_{\boldsymbol{\theta}}\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)=\nabla_{\boldsymbol{\theta}}\sum_{\boldsymbol{y}\in\Sigma^{*}}p_{\boldsymbol{\theta}}(\boldsymbol{y})\log\frac{p_{\boldsymbol{\theta}}(\boldsymbol{y})}{q(\boldsymbol{y})}.$ (18)
The standard approach to estimating the gradient is to apply the gradient operator inside the summation and then use the log-derivative trick to arrive at the following:
$\nabla_{\boldsymbol{\theta}}\textsc{kl}$ $=\sum_{\boldsymbol{y}\in\Sigma^{*}}\nabla_{\boldsymbol{\theta}}p_{\boldsymbol{\theta}}(\boldsymbol{y})\log\frac{p_{\boldsymbol{\theta}}(\boldsymbol{y})}{q(\boldsymbol{y})}+\nabla_{\boldsymbol{\theta}}p_{\boldsymbol{\theta}}(\boldsymbol{y})=\operatorname*{\mathbb{E}}\left[\left(\log\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y})}{q(\boldsymbol{Y})}+1\right)\nabla_{\boldsymbol{\theta}}\log p_{\boldsymbol{\theta}}(\boldsymbol{Y})\right].$ (19)
Since $\operatorname*{\mathbb{E}}[\nabla_{\boldsymbol{\theta}}\log p_{\boldsymbol{\theta}}(\boldsymbol{Y})]=0$, the gradient simplifes to
$\nabla_{\boldsymbol{\theta}}\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)=\operatorname*{\mathbb{E}}\left[\log\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y})}{q(\boldsymbol{Y})}\nabla_{\boldsymbol{\theta}}\log p_{\boldsymbol{\theta}}(\boldsymbol{Y})\right].$ (20)
Therefore, the Monte Carlo estimator of the gradient is
$\nabla\delta_{\textsc{mc}}=\frac{1}{M}\sum_{m=1}^{M}\log\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}\nabla_{\boldsymbol{\theta}}\log p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)}).$ (21)
Next, we derive the Rao–Blackwellized Monte Carlo estimator of the gradient. However, we first restate Theorem 2.2 in (Malagutti et al., 2024), which will prove useful.
Theorem 6 (Malagutti et al. (2024); Theorem 2.2).
Let $p_{\boldsymbol{\theta}}$ and $q$ be two language models over $\Sigma$. Furthermore, let $\vec{p}$ be the prefix probability function of $p$. Furthermore, we assume that $\mathrm{H}(p,q)<\infty$. Then, the following equality holds
where we treat $\vec{p}(\cdot\mid\boldsymbol{y})$ and $\vec{q}\left(\cdot\mid\boldsymbol{y}\right)$ as probability distributions over $\overline{\Sigma}$.
We refer the reader to Malagutti et al. (2024) for the proof. Next, to derive the Rao–Blackwellized Monte Carlo estimator of the gradient, we take the gradient of the local KL as stated in the following theorem.
Theorem 7 (Gradient of the local KL).
Let $p_{\boldsymbol{\theta}}$ and $q$ be two language models over $\Sigma$ and $\vec{p}_{\boldsymbol{\theta}}$ the prefix probability function of $p_{\boldsymbol{\theta}}$. Furthermore, we assume that $\mathrm{H}(p_{\boldsymbol{\theta}},q)<\infty$. Then, the following equality holds
$\nabla_{\boldsymbol{\theta}}\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)=\sum_{\boldsymbol{y}\in\Sigma^{*}}\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{y})\operatorname*{\mathbb{E}}_{\overline{Y}}\Big{[}\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}\mid\boldsymbol{y})}{\vec{q}(\overline{Y}\mid\boldsymbol{y})}\cdot\left(\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{Y}\mid\boldsymbol{y})+\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{y})\right)\Big{]}.$ (23)
Proof.
See App. F. ∎
We then construct the Monte Carlo estimator of the gradient using the above theorem, which naturally results in an unbiased Rao–Blackwellized Monte Carlo estimator of the gradient.
$\nabla\delta_{\textsc{rb}}=\frac{1}{M}\sum_{m=1}^{M}\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{Y}_{n}}\Big{[}\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}\cdot\left(\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}^{(m)}_{<n})+\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})\right)\Big{]}.$ (24)
The above estimator is unbiased, as it is the Monte Carlo estimator of the local KL gradient.
Theorem 8.
Let $p_{\boldsymbol{\theta}}$ and $q$ be two language models over $\Sigma$ and $\vec{p}_{\boldsymbol{\theta}}$ the prefix probability function of $p_{\boldsymbol{\theta}}$. Let $\nabla\delta_{\textsc{mc}}$ be the Monte Carlo estimator of the true gradient, denoted with $G\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\nabla_{\boldsymbol{\theta}}\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)$, as defined in Eq. 21 and $\nabla\delta_{\textsc{rb}}$ be the Rao–Blackwellized estimator of the gradient defined in Eq. 24. Assuming $\mathrm{H}(p_{\boldsymbol{\theta}},q)<\infty,\mathop{\mathrm{Var}}\left[{\normalcolor\nabla\delta_{\textsc{rb}}}\right]<\infty,\mathop{\mathrm{Var}}\left[{\normalcolor\nabla\delta_{\textsc{mc}}}\right]<\infty$, we have
$\operatorname*{\mathbb{E}}\left[\left\lVert\nabla\delta_{\textsc{rb}}-G\right\rVert^{2}\right]\leq\operatorname*{\mathbb{E}}\left[\left\lVert\nabla\delta_{\textsc{mc}}-G\right\rVert^{2}\right].$ (25)
Proof.
See § F.1. ∎
4.1 Off-policy Gradient
So far, we have discussed how to estimate the gradient of $\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)$ using samples drawn from the current policy $p_{\boldsymbol{\theta}}$. Importantly, the reason that we discussed the gradient of the estimators separately, rather than relying on the automatic differentiation, is that the samples are drawn from the $p_{\boldsymbol{\theta}}$, which depends on $\boldsymbol{\theta}$.
@@ 4.1 Off-policy Gradient
Although on-policy reinforcement learning algorithms sample from the same policy that is being optimized, in practice and for efficiency reasons, we often collect large batches of samples in parallel with the optimization loop. As a result, these samples may be generated from a slightly outdated version of the policy, denoted $p_{\boldsymbol{\theta}_{\text{old}}}$. To compute the KL divergence using samples from $p_{\boldsymbol{\theta}_{\text{old}}}$, we first write the KL as the expectation under $p_{\boldsymbol{\theta}_{\text{old}}}$ as
@@
Therefore, the Monte Carlo estimator using samples $\boldsymbol{Y}^{(1)},\cdots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p_{\boldsymbol{\theta}_{\text{old}}}$ is
$\delta_{\textsc{mc}}^{\text{old}}=\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y}^{(m)})}\log\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}.$ (27)
Given the unbiasedness proof of the Rao–Blackwellization in Thm. 5, we can similarly write
$\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)=\operatorname*{\mathbb{E}}_{\boldsymbol{Y}\sim p_{\boldsymbol{\theta}_{\text{old}}}}\Bigg{[}\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y})}{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y})}\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{Y}_{n}}\left[\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}_{<n})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}_{<n})}\right]\Bigg{]}.$ (28)
Therefore, the Rao–Blackwellized Monte Carlo estimator using samples $\overline{\boldsymbol{Y}}^{(1)},\cdots,\overline{\boldsymbol{Y}}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p_{\boldsymbol{\theta}_{\text{old}}}$ is
$\delta_{\textsc{rb}}^{\text{old}}=\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y}^{(m)})}\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{Y}_{n}}\left[\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}_{<n}^{(m)})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}_{<n}^{(m)})}\right].$ (29)
Since $\delta_{\textsc{mc}}^{\text{old}}$ and $\delta_{\textsc{rb}}^{\text{old}}$ use samples from the old policy that does not depend on $\boldsymbol{\theta}$, we can apply automatic differentiation to compute the estimate of the KL gradient by computing the gradient of $\delta_{\textsc{mc}}^{\text{old}}$ and $\delta_{\textsc{rb}}^{\text{old}}$.
Figure 1: Comparing the bias, variance, and consistency of the estimators as the sample size increases. The $\delta_{\textsc{cv}}$ estimator with $\alpha=1$ exhibits a higher standard deviation, particularly for neutral and negative prompts, where the variance becomes extremely large. In contrast, the Rao–Blackwellized estimator, $\delta_{\textsc{rb}}$, achieves the lowest standard deviation.
@@ 5 Experiments
We use the sentiment control task as the testbed to empirically evaluate our theoretical findings on the KL estimators. Concretely, the reference model, denoted as $q$, is the GPT-IMDB [^9] Specifically, we use https://huggingface.co/lvwerra/gpt2-imdb. model, i.e., a GPT-2 (Radford et al., 2019) model fine-tuned on imdb corpus (Maas et al., 2011). The goal of the task is to fine-tune this language model such that the samples from it are movie reviews with a positive sentiment. The fine-tuned language model is denoted with $p_{\boldsymbol{\theta}}$. In the following experiments, we estimate the KL divergence between $p_{\boldsymbol{\theta}}$ and $q$.
5.1 Understanding the KL Estimators
In this experiment, we empirically evaluate the bias, variance, and consistency of various KL estimators. To obtain $p_{\boldsymbol{\theta}}$, we fine-tune $q$ with direct preference optimization (DPO; Rafailov et al., 2023) on a sample of 5,000 data points from the IMDB training set. To create the preference data required for DPO training, following Rafailov et al. (2023), we sample $4$ responses for each prompt and create $6$ pairs per prompt. To determine the preferred response in each pair, we employ a binary sentiment classifier, [^10] Specifically, we use https://huggingface.co/lvwerra/distilbert-imdb. selecting the response with the higher probability of positive sentiment. Upon successful fine-tuning, $p_{\boldsymbol{\theta}}$ should assign a higher probability mass to movie reviews with positive sentiment while maintaining a low KL divergence with $q$. We then evaluate this KL divergence using our estimators to assess their reliability in measuring distributional shifts induced by fine-tuning.
We sample 512 examples from the evaluation set of the IMDB corpus. For each example, we randomly select a prefix length between 2 and 8 tokens and use it as the prompt. Then, for each prompt, we generate 4000 samples from the DPO fine-tuned model, $p_{\boldsymbol{\theta}}$. Using these samples, we compute the MC, CV, and RB estimators and empirically estimate their standard deviation. We further implement and compare the Horvitz–Thompson estimator; we refer the reader to App. A for a detailed discussion. We compute the Monte Carlo estimator with control variate, $\delta_{\textsc{cv}}$, twice: once using the optimal $\alpha$ value derived from 1000 samples and once with $\alpha=1$, matching the setup of Schulman (2020).

### v3

@@ 4 Estimating the Gradient
See App. D. ∎
We then construct the Monte Carlo estimator of the gradient using the above theorem, which naturally results in the following unbiased, Rao–Blackwellized Monte Carlo estimator of the gradient:
$\boldsymbol{\delta}_{\textsc{rb}}$ $=\frac{1}{M}\sum_{m=1}^{M}\sum_{n=1}^{|\boldsymbol{Y}^{(m)}|}\operatorname*{\mathbb{E}}_{Y}\Big[\log\frac{\vec{p}_{\boldsymbol{\theta}}(Y\mid\boldsymbol{Y}^{(m)}_{<n})}{\vec{q}(Y\mid\boldsymbol{Y}^{(m)}_{<n})}\cdot\left(\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)}_{<n})+\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(Y\mid\boldsymbol{Y}^{(m)}_{<n})\right)\Big$]. (16)
The above estimator is unbiased, as it is the MC estimator of Eq. 15.
Theorem 5.
Assuming $\mathop{\mathrm{Var}}\left[{\normalcolor\boldsymbol{\delta}_{\textsc{rb}}}\right]<\infty,\mathop{\mathrm{Var}}\left[{\normalcolor\boldsymbol{\delta}_{\textsc{mc}}}\right]<\infty$, we have
$\operatorname*{\mathbb{E}}\left[\left\lVert\boldsymbol{\delta}_{\textsc{rb}}-\boldsymbol{G}\right\rVert^{2}\right]\leq\operatorname*{\mathbb{E}}\left[\left\lVert\boldsymbol{\delta}_{\textsc{mc}}-\boldsymbol{G}\right\rVert^{2}\right].$ (17)
@@
See § D.1. ∎
Off-policy Gradient.
@@
So far, we have discussed how to estimate the gradient of $\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)$ using samples drawn from the current policy $p_{\boldsymbol{\theta}}$. Crucially, we derive the gradient manually rather than relying on automatic differentiation because the samples depend on $\boldsymbol{\theta}$ through $p_{\boldsymbol{\theta}}$. However, in practice and for efficiency reasons, we often collect large batches of samples in parallel with the optimization loop. As a result, these samples are generated from a slightly outdated version of the policy, denoted $p_{\boldsymbol{\theta}_{\text{old}}}$. To compute the KL divergence using samples from $p_{\boldsymbol{\theta}_{\text{old}}}$, we first write the KL as the expectation under $p_{\boldsymbol{\theta}_{\text{old}}}$ as
@@
Therefore, the MC estimator using samples $\boldsymbol{Y}^{(1)},\ldots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p_{\boldsymbol{\theta}_{\text{old}}}$ is
$\mu_{\textsc{mc}}^{\text{old}}=\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y}^{(m)})}\log\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}.$ (19)
Given the unbiasedness proof of the Rao–Blackwellization in Thm. 2, we can similarly write
$\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)=\operatorname*{\mathbb{E}}_{\boldsymbol{Y}\sim p_{\boldsymbol{\theta}_{\text{old}}}}\Bigg[\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y})}{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y})}\sum_{n=1}^{|\boldsymbol{Y}^{(m)}|}\operatorname*{\mathbb{E}}_{Y}\left[\log\frac{\vec{p}_{\boldsymbol{\theta}}(Y\mid\boldsymbol{Y}_{<n})}{\vec{q}(Y\mid\boldsymbol{Y}_{<n})}\right]\Bigg].$ (20)
Therefore, the Rao–Blackwellized MC estimator using samples $\boldsymbol{Y}^{(1)},\ldots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p_{\boldsymbol{\theta}_{\text{old}}}$ is
$\mu_{\textsc{rb}}^{\text{old}}=\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{p_{\boldsymbol{\theta}_{\text{old}}}(\boldsymbol{Y}^{(m)})}\sum_{n=1}^{|\boldsymbol{Y}^{(m)}|}\operatorname*{\mathbb{E}}_{Y}\left[\log\frac{\vec{p}_{\boldsymbol{\theta}}(Y\mid\boldsymbol{Y}_{<n}^{(m)})}{\vec{q}(Y\mid\boldsymbol{Y}_{<n}^{(m)})}\right].$ (21)
Since $\mu_{\textsc{mc}}^{\text{old}}$ and $\mu_{\textsc{rb}}^{\text{old}}$ use samples from the old policy that does not depend on $\boldsymbol{\theta}$, we can apply automatic differentiation to compute the estimate of the KL gradient by computing the gradient of $\mu_{\textsc{mc}}^{\text{old}}$ and $\mu_{\textsc{rb}}^{\text{old}}$.
@@ 5 Experiments
We use the sentiment control task as the testbed to empirically evaluate our theoretical findings on the KL estimators. Concretely, the reference model, denoted as $q$, is the GPT-IMDB [^14] Specifically, we use https://huggingface.co/lvwerra/gpt2-imdb. model, i.e., a GPT-2 [28] model fine-tuned on imdb corpus [23]. The goal of the task is to fine-tune this language model such that the samples from it are movie reviews with a positive sentiment. The fine-tuned language model is denoted with $p_{\boldsymbol{\theta}}$. In the following experiments, we estimate the KL divergence between $p_{\boldsymbol{\theta}}$ and $q$. We provide a code snippet for implementing the RB estimator in § F.1.
5.1 Analyzing the KL Estimators
In this experiment, we empirically evaluate the bias, variance, and consistency of various KL estimators. To obtain $p_{\boldsymbol{\theta}}$, we fine-tune $q$ with direct preference optimization [DPO; 29] on a sample of 5,000 data points from the IMDB training set. To create the preference data required for DPO training, following Rafailov et al. [29], we sample 44 responses for each prompt and create 66 pairs per prompt.
Table 1: Estimated value $\pm$ empirical standard deviation of different estimators. When aggregating over prompts, $\mu_{\textsc{ht}}$ and $\mu_{\textsc{cv}}$ fail to significantly reduce the variance of $\mu_{\textsc{mc}}$. RB estimator, however, achieves the lowest standard deviation.